# Ano : Faster is Better in Noisy Landscapes

### v1

@@ D.3 Full Normalized Score for PPO
Baseline
Adam $7656.00\pm 812.20$ 91.19 $\mathbf{-0.91}\pm\mathbf{0.34}$ 100 $733.92\pm 107.34$ 90.99 $\mathbf{3516.64}\pm\mathbf{475.14}$ 100 $4322.00\pm 254.93$ 79.83 92.40
Adan $6408.00\pm 1525.14$ 76.32 $0.98\pm 0.21$ 99.58 $633.80\pm 44.93$ 78.58 $2033.76\pm 85.42$ 57.83 $2100.60\pm 951.20$ 38.80 70.22
Lion $1420.00\pm 256.88$ 16.91 $1.96\pm 1.50$ 93.84 $564.44\pm 195.39$ 69.98 $2477.12\pm 490.65$ 70.44 $3680.80\pm 1057.34$ 67.99 63.83
Grams $\mathbf{8396.00}\pm\mathbf{1565.13}$ 100 $1.07\pm 0.55$ 99.09 $644.64\pm 48.73$ 79.92 $2133.76\pm 234.64$ 60.68 $3646.70\pm 1139.41$ 67.36 81.41
Ano (Ours) $8280.00\pm 916.31$ 98.62 $1.03\pm 0.21$ 99.30 $\mathbf{806.56}\pm\mathbf{114.49}$ 100 $2808.68\pm 465.32$ 79.87 $5122.12\pm 783.72$ 94.61 94.48
Anolog (Ours) $7620.00\pm 1225.14$ 90.76 $1.07\pm 0.26$ 99.08 $743.00\pm 191.30$ 92.12 $2815.50\pm 430.55$ 80.06 $\mathbf{5413.75}\pm\mathbf{1873.65}$ 100 92.40
@@
Adam $7880.00\pm 1218.48$ 82.46 $\mathbf{-0.93}\pm\mathbf{0.15}$ 100 $663.87\pm 178.56$ 90.45 $2964.16\pm 601.40$ 98.14 $4222.90\pm 256.83$ 81.50 90.51
Adan $4672.00\pm 2542.63$ 48.89 $1.08\pm 0.27$ 99.11 $693.96\pm 87.13$ 94.55 $\mathbf{3020.48}\pm\mathbf{419.22}$ 100 $4540.20\pm 996.34$ 87.62 86.03
Lion $1296.00\pm 164.16$ 13.56 $2.73\pm 1.92$ 89.46 $550.96\pm 134.29$ 75.07 $2554.80\pm 546.74$ 84.58 $3267.00\pm 915.37$ 63.05 65.14
Grams $7536.00\pm 240.38$ 78.86 $1.41\pm 0.46$ 97.19 $704.68\pm 123.61$ 96.01 $1864.92\pm 204.72$ 61.74 $4962.38\pm 1373.46$ 95.77 85.91
Ano (Ours) $\mathbf{9556.00}\pm\mathbf{2370.99}$ 100 $1.13\pm 0.43$ 98.83 $\mathbf{733.95}\pm\mathbf{235.51}$ 100 $2862.24\pm 365.46$ 94.76 $\mathbf{5181.70}\pm\mathbf{1210.60}$ 100 98.72
Table 9: Comparison of the average performance (± IC95%) and normalized scores of different optimizers across Atari environments.
@@ E.1 Optimizers Hyperparameters
Adan 0.92 0.92 0.99 1e-3 1e-2
Ano 0.95 0.995 1e-4 1e-2
Lion 0.9 0.99 1e-4 1e-2
Grams 0.9 0.999 1e-3 1e-2
@@
BERT (base) AdamW 0.9 0.99 7e-5 1e-2
@@
Ano 0.92 0.999 2e-5 1e-2
Lion 0.9 0.99 7e-6 1e-2
Grams 0.9 0.999 7e-5 1e-2
Anolog 0.999 2e-5 1e-2
@@
Grams 0.9 0.95 3e-3

### v2

@@ D.2 Full Normalized Score for PPO
Default
Adam $7615.00\pm 1299.89$ 94.07 $1.08\pm 0.21$ 97.62 $665.35\pm 64.89$ 78.71 $3443.30\pm 240.68$ 100 $4257.80\pm 135.90$ 88.18 91.71
RMSprop $7686.67\pm 859.43$ 94.96 $0.67\pm 0.22$ 100 $798.00\pm 118.66$ 94.40 $3031.13\pm 410.22$ 88.03 $4585.67\pm 538.44$ 94.96 94.47
Adan $6480.00\pm 702.14$ 80.05 $0.91\pm 0.23$ 98.58 $638.35\pm 18.08$ 75.51 $2106.90\pm 110.74$ 61.19 $2665.00\pm 708.69$ 55.19 74.11
Lion $1392.00\pm 139.66$ 17.20 $1.46\pm 0.65$ 95.42 $508.15\pm 91.28$ 60.11 $2432.35\pm 234.61$ 70.64 $3768.00\pm 520.95$ 78.03 64.28
Grams $7908.00\pm 687.61$ 97.69 $0.93\pm 0.24$ 98.48 $633.80\pm 23.52$ 74.97 $2234.40\pm 130.73$ 64.89 $3670.12\pm 634.75$ 76.00 82.41
Ano (Ours) $8095.00\pm 494.70$ 100 $0.97\pm 0.14$ 98.26 $845.35\pm 56.39$ 100 $2813.00\pm 243.04$ 81.69 $4828.80\pm 386.73$ 100 95.99
Anolog (Ours) $7485.00\pm 1010.66$ 92.46 $0.98\pm 0.14$ 98.19 $751.05\pm 74.60$ 88.84 $2983.00\pm 236.42$ 86.63 $4773.75\pm 602.39$ 98.86 93.00
@@
Adam $6430.00\pm 864.51$ 74.55 $0.98\pm 0.20$ 99.14 $549.75\pm 51.42$ 66.39 $406.90\pm 105.28$ 14.40 $4486.62\pm 683.06$ 75.27 65.95
[…]
Adan $4840.00\pm 2601.35$ 56.12 $0.95\pm 0.23$ 99.30 $754.20\pm 51.44$ 91.08 $2647.20\pm 534.22$ 93.71 $4524.75\pm 448.74$ 75.91 83.22
Lion $1324.00\pm 218.94$ 15.35 $2.38\pm 1.03$ 90.98 $574.55\pm 73.33$ 69.38 $2232.85\pm 364.56$ 79.04 $3759.75\pm 808.26$ 63.07 63.57
Grams $7715.00\pm 627.92$ 89.45 $1.35\pm 0.75$ 96.95 $690.40\pm 76.27$ 83.37 $1989.15\pm 201.64$ 70.42 $5049.25\pm 624.43$ 84.71 84.98
Ano (Ours) $8625.00\pm 1870.44$ 100 $0.91\pm 0.21$ 99.55 $828.10\pm 67.66$ 100 $2824.85\pm 226.30$ 100 $5960.88\pm 912.36$ 100 99.91
Anolog (Ours) $1470.00\pm 1176.65$ 17.04 $1.06\pm 1.06$ 98.67 $543.60\pm 111.65$ 65.64 $832.50\pm 125.46$ 29.47 $1323.50\pm 1694.70$ 22.20 46.61
Best Version
Adam $7615.00\pm 1299.89$ 88.29 $1.08\pm 0.21$ 97.62 $665.35\pm 64.89$ 80.35 $3443.30\pm 240.68$ 100 $4257.80\pm 135.90$ 71.43 87.54
RMSprop $7686.67\pm 859.43$ 89.12 $0.67\pm 0.22$ 100 $798.00\pm 118.66$ 96.37 $3031.13\pm 410.22$ 88.03 $4585.67\pm 538.44$ 76.93 90.09
Adan $4840.00\pm 2601.35$ 56.12 $0.95\pm 0.23$ 98.35 $754.20\pm 51.44$ 91.08 $2647.20\pm 534.22$ 76.88 $4524.75\pm 448.74$ 75.91 79.67
Lion $1392.00\pm 139.66$ 16.14 $1.46\pm 0.65$ 95.42 $508.15\pm 91.28$ 61.36 $2432.35\pm 234.61$ 70.64 $3768.00\pm 520.95$ 63.21 61.36
Grams $7715.00\pm 627.92$ 89.45 $1.35\pm 0.75$ 96.03 $690.40\pm 76.27$ 83.37 $1989.15\pm 201.64$ 57.77 $5049.25\pm 624.43$ 84.71 82.26
Ano (Ours) $8625.00\pm 1870.44$ 100 $0.91\pm 0.21$ 98.60 $828.10\pm 67.66$ 100 $2824.85\pm 226.30$ 82.04 $5960.88\pm 912.36$ 100 96.13
Anolog (Ours) $7485.00\pm 1010.66$ 86.78 $0.98\pm 0.14$ 98.19 $751.05\pm 74.60$ 90.70 $2983.00\pm 236.42$ 86.63 $4773.75\pm 602.39$ 80.08 88.48
Table 9: Comparison of the average performance ($\pm$ IC95%) and normalized scores of different optimizers across Atari environments.
@@ E.1 Optimizers Hyperparameters
Adan 0.95 0.92 0.9 1e-3 1e-2
Ano 0.95 0.95 1e-4 1e-2
Lion 0.92 0.99 1e-4 1e-2
Grams 0.9 0.99 1e-3 1e-2
@@
BERT (base) AdamW 0.95 0.99 7e-5 1e-2
@@
Ano 0.9 0.95 2e-5 1e-2
Lion 0.9 0.9 7e-6 1e-2
Grams 0.92 0.999 7e-5 1e-2
Anolog 0.95 2e-5 1e-2
@@
Grams 0.9 0.999 3e-3
RMSprop 0.999 3e-3
Anolog 0.99 3e-3
Best Version (SAC)
SAC Adam 0.9 0.999 3e-4
Adan 0.97 0.92 0.96 3e-3
Ano 0.92 0.99 3e-4
Lion 0.92 0.99 3e-4
Grams 0.9 0.999 3e-3
RMSprop 0.99 3e-4
Anolog 0.999 3e-4
Best Version (PPO)
PPO Adam 0.9 0.999 3e-4
Adan 0.97 0.92 0.96 3e-3
Ano 0.95 0.99 3e-4
Lion 0.92 0.99 3e-4
Grams 0.9 0.999 3e-3
RMSprop 0.99 3e-4